### v4

# Vertex-Minimal Paper Tori

Abstract
A paper torus is an embedded polyhedral torus that is isometric to a flat torus in the intrinsic sense. We prove that there does not exist a paper torus with $7$ vertices, and that there does exist a paper torus with $8$ vertices.
1 Introduction
1.1 Context and Main Results
A flat torus is a quotient of the form $\mbox{\boldmath{$R$}}^{2}/\Lambda$, where $\Lambda$ is a lattice of translations of $\mbox{\boldmath{$R$}}^{2}$. More concretely, a flat torus is what you get when you identify the opposite sides of a parallelogram by translations.
A polyhedral torus is a continuous, piecewise affine embedding $\phi:T\to\Omega\subset\mbox{\boldmath{$R$}}^{3}.$ Here $T$ is a flat torus that has been triangulated, and $\phi$ is a continuous embedding that is affine on each triangle of the triangulation. If, additionally, $\phi$ is an affine isometry on each triangle, we say that $\phi$ is flat. When $\phi$ is a flat embedding, $\phi$ tells how to build a torus in $\mbox{\boldmath{$R$}}^{3}$ out of finitely many triangles so that the cone angle around each vertex is $2\pi$. We call a flat embedded polyhedral torus a paper torus. Some authors would call this an origami torus.
Surprisingly, paper tori exist. The 1960 paper of Y. Burago and V. Zalgaller [BZ1] gives the first constructions. The 1985 paper [BZ2] proves that one can realize every isometry class of flat torus as a paper torus. Their construction produces examples with thousands of faces. The 2024 preprint of F. Lazarus and F. Tallerie [LT] gives a universal combinatorial type of triangulation which does the job simultaneously for all isometry types.
The works of P. Arnoux, S. Lelievre, and A. Malaga [ALM] and (independently) T. Tsuboi [T] give a different construction which achieves every isometry class of flat torus as a paper torus, using far fewer faces than in the other cited works. With minor differences in the final results, both papers essentially prove that the paper tori described by U. Brehm [Br] in 1978 achieve every isometry class of flat torus. In 2025, Vincent Tugayé [Tu], discovered a $9$-vertex paper torus.
One needs at least $7$ vertices to make a paper torus because there is no $6$ vertex triangulation of a torus. There is a $7$-vertex triangulation of a torus, and it is unique up to combinatorial isomorphism. This triangulation, sometimes called the Moebius triangulation, is shown in Figure 1.1 below. The Moebius triangulation gives the famous embedding of the complete graph $K_{7}$ in a torus. Figure 1.1 shows part of the universal cover of the Moebius torus, as well as a fundamental domain for it.
Figure 1.1: Part of the universal cover of the Moebius triangulation
In $1949$, Á. Császár [Cs] showed that the Moebius torus has a polyhedral embedding. However, this embedding is not flat. See e.g. [G], [HLZ], and [L] for more detail and a discussion of related topics. The 2019 Ph.D. thesis of P. Quintanar Cortés, [QC] makes significant progress towards showing that there is no $7$-vertex paper torus.
The results above leave open the question as to the minimum number of vertices needed to make a paper torus. The results in this paper settle the question.
Theorem 1.1
There does not exist a $7$-vertex paper torus.
Theorem 1.2
There exists an $8$-vertex paper torus.
The example in Theorem 1.2 has all $8$ points in its convex hull boundary and $2$-fold symmetry. We will discuss it below in great detail.
1.2 Seven Vertices: Proof Sketch
The proof of Theorem 1.1 is actually rather easy. Here is the key step.
Lemma 1.3 (Hull)
Suppose $\Omega$ is an embedded $7$-vertex polyhedral torus, not necessarily flat. Let $H$ be the convex hull of $\Omega$. Suppose $\Omega$ has all $7$ vertices in $\partial H$. Then $\Omega$ has at least one vertex $P$ such that all $6$ triangles of $\Omega$ incident to $P$ are in $\partial H$.
I will deduce from the Hull Lemma that a paper torus must have a vertex $P$ contained in the interior of the convex hull. I will then use Crofton’s formula to show that the cone angle at $P$ exceeds $2\pi$. This contradiction finishes the proof of Theorem 1.1.
The Hull Lemma is a consequence of a stronger result contained in the 1991 paper [BE] of J. Bokowski and A. Eggert. Bokowski and Eggert use oriented matroids to classify the combinatorial types of $7$-vertex polyhedral tori. One consequence of their classification is the Hull Theorem: there are no embedded $7$-vertex polyhedral tori at all having all $7$ vertices on their convex hull boundary. (See the remarks after [BE, Theorem 3.7]. This is my name for their result.) In my opinion, had these authors asked the flatness question for $7$-vertex polyhedral tori, they most likely would have been able to answer it.
In order to keep this paper self-contained, I will give a light and entirely combinatorial computer-assisted proof of the Hull Lemma. I will also derive the Hull Theorem from the Hull Lemma, though this is not needed for our results.
1.3 Eight Vertices: Experimentation
Motivated by our proof of Theorem 1.1, I tried to rule out the existence of $8$-vertex paper tori having all $8$ vertices on their convex hull boundary. This did not work, so I decided to find these kinds of examples and study them.
There are $7$ combinatorial types of triangulation of the torus having $8$ vertices. Of these $7$ there is (in my opinion) a worst triangulation and a best triangulation. The worst triangulation is obtained from the Moebius triangulation by subdividing one of the triangles. The best triangulation has degree sequence $66666666$ and is vertex-transitive. This triangulation is very much like the Moebius triangulation. Figure 1.2 below shows part of the universal cover of this triangulation, as well as a fundamental domain. I will explain the meaning of the blue triangles below.
For all but the worst triangulation, I generated lots of embedded examples by randomly sampling collections of $8$ points on the unit sphere and testing the embedding condition.
Figure 1.2: The universal cover of the best $8$ vertex triangulation
Next, I ran a hill-climbing algorithm which started with random embedded examples on the unit sphere and repeatedly perturbed them – not necessarily keeping them on the unit sphere – with the goal of minimizing an objective flatness function. The flatness function computes the maximum deviation of a cone angle from being $2\pi$. After a lot of experimentation, I implemented $4$ tricks that helped immensely.
  1. 1.
    It was useful to first perturb the points within an ellipsoid surface that contains them, and then to perturb the ellipsoid. This method is guaranteed to keep all vertices on the convex hull boundary.
  2. 2.
    It was extremely useful to control the combinatorics on the convex hull boundary. The convex hull boundary has $12$ faces, some of which – say $K$ of them – are also faces in the triangulation of the torus. Call $K$ the face number. I noticed that the examples with large face number never led to flat examples. So I eventually constrained my search to examples with low face number.
  3. 3.
    It helped to bound some auxiliary quantities, like the minimum angles and dihedral angles, to prevent degenerations. Bounds of about $0.00001$ on the dihedral angles worked well.
  4. 4.
    It helped to force the examples to have $2$-fold rotational symmetry. This simplification came later. After I found some asymmetric examples, Samuel Lelievre asked me about symmetric examples and I tweaked the program. The program works more reliably after incorporating the symmetry.
1.4 Meet the Pup Tent
With all the tricks above in place, and a lot of playing around, I found some embedded and $2$-fold symmetric examples of face number $6$ which are flat up to $10^{-16}$. I took the nicest one and simplified the coordinates, feeding it back into the program to make it near-flat again. Then I implemented a high precision version of Newton’s method and improved my example to one that was flat up to $10^{-400}$. Using a different method, Peter Doyle computed the first million digits of my example!
Here are the coordinates for my example, truncated after $32$ digits. Theorem 1.2 asserts more specifically that this example is within $10^{-32}$ of an $8$-vertex paper torus.
[…]
The 0th row of this array gives coordinates for vertex 0, and so on. The example has $2$-fold symmetry which reflects in the $Z$-axis and implements the vertex permutation $(04)(13)(26)(57)$. You can operate my software and generate up to about $1000$ digits of precision.
Based on his computer plots, Doyle suggested that I call this example the pup tent. Let us say more formally that a pup tent is an $8$-vertex paper torus with $2$-fold rotational symmetry, based on the best triangulation, and having $6$ faces in the convex hull boundary arranged combinatorially like the blue triangles in Figure 1.2.
Figure 1.3: The intrinsic structure.
Figure 1.3 shows a numerical computation of the intrinsic geometric structure of the faces of the pup tent. We are showing part of the universal cover of the triangulation. You can generate your own version of Figure 1.3 using the following $8$ vertices and the two generators $\lambda_{1},\lambda_{2}$ of the deck group.
[…]
Figure 1.3 looks quite tame, but the extrinsic geometry of the pup tent is beautiful and very intricate. Figure 1.4 shows one of the slices of the pup tent by a plane parallel to the $Z$ axis. The solid handlebody bounded by the torus is colored orange. This slice is very spiky and sort of “barely embedded”, but the embeddedness is robust compared to the plot precision.
Figure 1.4: A horizontal slice of the pup tent
Figure 1.5 shows the slice by the $XZ$ plane, which goes through the axis of symmetry of the torus. The slice of the handlebody is an annulus. The two “parallel vertical segments” on top are actually each unions of two segments which meet in the middle at angle that is close to $\pi-.01$. We show a closeup of the top central area to highlight the fact that the two polygons really are disjoint.
Figure 1.5: The $XZ$ planar slice of the pup tent.
Figure 1.6 shows the slice by the $YZ$ plane, with some closeups.
Figure 1.6: The $YZ$ planar slice of the pup tent
The left side of Figure 1.7 shows a random slice of the pup tent.
Figure 1.7: Extrinsic and Intrinsic structures compared
Once we fix this direction we get a level set foliation of the pup tent by curves that lie in planes perpendicular to this direction. On the right side of Figure 1.7, we have drawn the level set foliation in the intrinsic flat structure and then lifted the picture to the universal cover. The highlighted loop on the right corresponds to the loop shown on the left.
In §6, I include Doyle’s plots of the pup tent. I also include a sheet of paper I made, after consulting with Alba Malaga and Noah Mongtomery, which you can fold up into your own example.
1.5 Eight Vertices: Proof Sketch
If we leave the variables $z_{0},z_{1},z_{2}$ unspecified, the example in Equation 1 is part of a $3$-parameter family of examples we call $T(z_{0},z_{1},z_{2})$. Let $\theta_{k}(z_{0},z_{1},z_{2})$ be the cone angle of $T(z_{0},z_{1},z_{2})$ at vertex $k$. Let
$F(z_{0},z_{1},z_{2})=(\theta_{0},\theta_{1},\theta_{2}).$
If $F(z_{0},z_{1},z_{2})=(2\pi,2\pi,2\pi)$ and $T(z_{0},z_{1},z_{2})$ is embedded then we have our proof. We use the Inverse Function Theorem. Let $dF$ be the Jacobian of $F$, computed at our example. We display $dF$ and its inverse up to $2$ digits:
$\left[\begin{matrix}-0.91\ldots&+0.74\ldots&+0.39\ldots\\ +0.74\ldots&-1.92\ldots&+1.14\ldots\\ +0.39\ldots&+1.14\ldots&-0.06\ldots\end{matrix}\right]\qquad\left[\begin{matrix}-0.56\ldots&+0.23\ldots&+0.76\ldots\\ +0.23\ldots&-0.04\ldots&+0.64\ldots\\ +0.76\ldots&+0.64\ldots&+0.57\ldots\end{matrix}\right]$ (3)
Notice that $dF$ is very well conditioned because $dF^{-1}$ is small. Our proof is an interplay of three things. First, we show that our example is robustly embedded in the sense that moderate perturbations keep it embedded. Second, we show that our example is very flat. Third we show, by bounding derivatives, that our Jacobian matrix stays fairly constant and well conditioned through a healthy neighborhood of our example. We then use an effective version of the Inverse Function Theorem so show that a true paper torus is quite near our example. In many parts of the proof, we work with a scale-up version of our example so that we can work with integer calculations.
Peter Doyle tells me that he very probably can show that my example is near a flat example by using the method in Matthew Ellison’s 2023 preprint [E].
Here is a quick corollary of our method.
Corollary 1.4
There exists a $6$-dimensional open manifold of pup tents.
Proof: We can normalize the vertices we work with so that the line of symmetry is the $Z$-axis. We can then perturb our points symmetrically by changing the $X$ and $Y$ coordinates but not the $Z$-coordinates. As long as the perturbation is small enough, the Inverse Function Theorem selects $Z$-coordinates for us which give a pup tent. We have $8$ free parameters to play with and then these are grouped into a $2$-dimensional foliation under the action of similarities of the $XY$-plane which fix the origin. A cross-section of this foliation gives us our $6$-dimensional open manifold of inequivalent pup tents. $\spadesuit$
1.6 Paper Organization
  • In §2, we will prove Theorem 1.1. At the end of §2 I derive the Hull Theorem from the Hull Lemma. This is not needed for any of the results in the paper.
  • In §3, we prove that our example $T$ from Equation 1 is $10^{-4}$-robustly embedded, meaning that if we change any of $z_{0},z_{1},z_{2}$ by less than $10^{-4}$ then the result is still embedded.
  • In §4, we prove that $T$ is $10^{-30}$-flat, meaning that the deviation of each cone angle from $2\pi$ is less than $10^{-30}$
  • In §5, we use the results from §3-4 and finish the proof of Theorem 1.2.
  • in §6, I include some additional pictures and some other information.
1.7 Computer Assistance
My proof is heavily computer assisted, though all the important calculations are done with exact integer arithmetic. The reader can download the code for Theorems 1.1 and 1.2 respectively at
[…]
I used ChatGPT-4o and ChatGPT-5 in various ways to help me with this project. Since this is my first time doing a project with the assistance of ChatGPT, I think that this warrants some discussion. I used ChatGPT to discuss high level mathematical ideas, to find typos and other inconsistencies in my paper, to format data both for the paper and for my computer code, and also to document the efficiency of my programs. I did not let ChatGPT compose prose for me and I ignored its writing advice. I want to retain $100$ percent of my own voice when I write something. Overall, I found ChatGPT to be a helpful and enthusiastic companion as I worked.
1.8 Acknowledgements
I thank Samuel Lelievre and Alba Malaga for telling me all about flat tori (over a period of some years) and in particular telling me about the flatness problem and supplying me with some historical context. The main thing that inspired me to work on this project was the infectious enthusiasm of Samuel and Alba. I also thank Peter Doyle, Fabian Lander, Noah Montgomery, Stepan Paul, and Saul Schleimer for helpful discussions about this material. I thank Frank Lutz for making his manifolds webpage, which lets you download the combinatorial data for triangulations of surfaces and other manifolds. This page was very useful to me.
I thank the IHES, the Hamilton Institute, and the Isaac Newton Institute, where I started working on this paper. This work is also supported by a Simons Sabbatical Fellowship and the Mazur Chair at IHES.
2 Proof of Theorem 1.1
2.1 Crofton’s Formula
Crofton’s Formula is a classic result from integral geometry. See [S]. It applies to any rectifiable arc on the unit sphere $S^{2}$. We just need this result for finite unions of arcs of great circles, which we call spherical polygons.
The space of oriented great circles in $S^{2}$ is canonically bijective with $S^{2}$ itself, and inherits a canonical probability measure $\sigma$. Given a spherical polygon $\gamma$ and a great circle $C$, we let $(C\cap\gamma)$ denote the number of intersection points. (We can ignore the finitely many great circles for which this is infinite.) Crofton’s formula is as follows.
${\rm length\/}(\gamma)=\pi\int_{\cal S}\#(C\cap\gamma)\ d\sigma.$ (4)
At least for spherical polygons, this has a swift proof. Both sides of Equation 4 are additive, so it suffices to prove the equation for a spherical polygon that is just an arc of a single great circle. By continuity, it suffices to prove the result for great circle arcs whose length is a rational multiple of $2\pi$. By additivity once again, it suffices to prove Crofton’s formula for the great circle itself. But then the formula is obvious.
Here is a well-known consequence of Crofton’s formula.
Lemma 2.1
Suppose $\gamma$ is a spherical polygon which is also a topological loop. If $\gamma$ has length at most $2\pi$ then $\gamma$ lies in some hemisphere of $S^{2}$.
Proof: Crofton’s formula says that some great circle $C$ intersects $\gamma$ at most once. But then $\gamma$ lies in one of the two hemispheres defined by $C$. When $\gamma$ has length exactly $2\pi$ we can shorten $\gamma$ a bit by cutting a small corner off. The shortened curve then lies in a hemisphere. Taking a limit as the cut corner tends to 0 in length, we see that $\gamma$ itself lies in a hemisphere. $\spadesuit$
2.2 Proof of the Main Theorem
Now suppose that $\Omega$ is a flat embedded $7$-vertex polyhedral torus. Let $H$ be the convex hull of $\Omega$. Call a vertex of $\Omega$ interior if it does not lie in $\partial H$. Given any vertex $P$ of $\Omega$ let $\Omega_{P}$ denote the union of $6$ triangles of $\Omega$ incident to $P$.
Lemma 2.2
$\Omega$ has an interior vertex.
Proof: Suppose that $\Omega$ does not have an interior vertex. By the Hull Lemma, $\Omega$ has a vertex $P$ such that all $6$ triangles of $\Omega$ incident to $P$ lie in $\partial H$. Let $\Omega_{P}$ be the union of triangles incident to $P$ as above.
A support plane for $H$ is a plane which intersects the boundary $\partial H$ and is disjoint from the interior of $H$. Since the cone angle of $\Omega$ at $P$ is $2\pi$, and since $\Omega_{P}\subset\partial H$, we see that $H$ cannot be strictly convex at $P$. Hence some support plane through $P$ intersects $\Omega_{P}$ in (at least) a line segment through $P$ that contains $P$ in its relative interior. But then the intersection of $\partial H$ with a small sphere centered at $P$ is a spherical polygon consisting of two arcs which connect the same pair of antipodal points.
Since the total length of $\gamma$ is $2\pi$, and since a great semicircle is the unique path on $S^{2}$ of length at most $\pi$ connecting two antipodal points, we see that $\gamma$ must be a union of two great semicircles. Hence $\Omega_{P}$ is contained in the union of two planes whose intersection contains $P$.
One of these two planes, $\Pi$, contains at least $3$ consecutive triangles of $\Omega_{P}$. But then $\Pi$ contains at least $5$ vertices of $\Omega$. Since the $1$-skeleton of the triangulation of $\Omega$ is the complete graph $K_{7}$, we see that $\Pi$ contains an embedded copy of $K_{5}$. This contradicts the fact that $K_{5}$ is not planar. $\spadesuit$
Now we know that our flat embedded $7$-vertex torus $\Omega$ has an interior vertex $P$.
Lemma 2.3
The cone angle $\theta$ at $P$ exceeds $2\pi$.
Proof: We translate so that $P$ is the origin. Let $S^{2}$ be the unit sphere. Let $\theta_{P}$ denote the cone angle of $\Omega$ at $P$. Let $\ell(\cdot)$ stand for length.
Let $\Omega_{P}$ be the union of $6$ triangles of $\Omega$ incident to $P$, as above. Let $\widehat{\Omega}_{P}$ denote the union of rays emanating from $P$ whose initial portions are contained in $\Omega_{P}$. Let $L_{P}=\widehat{\Omega}_{P}\cap S^{2}$. We have $\theta_{P}=\ell(L_{P})$.
Since $P$ is in the interior of the convex hull of $\Omega$ and since $\widehat{\Omega}_{P}$ contains all the vertices of this convex hull, we see that $\widehat{\Omega}_{P}$ cannot lie in any halfspace bounded by a plane through the origin. This means that $L_{P}$ cannot lie in any hemisphere of $S^{2}$. By Lemma 2.1, we have $\ell(L_{P})>2\pi$. Hence $\theta_{P}>2\pi$. $\spadesuit$
This completes the proof of the Main Theorem modulo the Hull Lemma.
2.3 Proof of the Hull Lemma
We suppose that $\Omega$ is an embedded $7$-vertex polyhedral torus having $7$ points on its convex hull $H$. We can perturb so that the points are in general position. This makes $H$ into a triangulated solid polyhedron with $7$ vertices, $15$ edges, and $10$ faces. One property we will use repeatedly is that $\Omega$ is neighborly: Every two vertices of $\Omega$ are joined by an edge in the triangulation. In particular, every edge of $H$ is an edge of the triangulation of $\Omega$.
We call the $15$ edges in $\partial H$ the external edges. We call the remaining $6$ edges internal edges. We say that an internal edge pattern is a choice of $6$ distinguished edges from the $1$-skeleton of the triangulation, normalized (by symmetry) so that the first internal edge is $(01)$. There are $\binom{20}{5}=15504$ different internal edge patterns.
We say that an external triangle of $\Omega$ is one that lies in $\partial H$, and an internal triangle is one that does not. Each internal edge is incident to two internal triangles.
Lemma 2.4
An internal triangle cannot be bounded by $3$ external edges.
Proof: Suppose that such a triangle exists. Call it $\tau$. If all three edges of $\tau$ lie in $\partial H$ and $\tau\not\subset\partial H$ then $\tau$ separates $H$ into two components, both of which contain vertices of $\Omega$ in their interior. Any path in $H$ connecting two such vertices must intersect $\tau$. On the other hand $\Omega-\tau$ is path connected. This is a contradiction. $\spadesuit$
By Lemma 2.4, the internal edge pattern determines the set of internal triangles and the set of external triangles. It is worth pointing out a triangle of $\partial H$ is not necessarily a triangle of $\Omega$.
Since $\Omega$ is neighborly, the degree of a vertex in the triangulation of $\partial H$ is just the number of external edges incident to that vertex, and this is the same as $6$ minus the number of internal edges incident to the vertex. In particular, since $H$ is a convex polyhedron and $\partial H$ is triangulated, the degree of each vertex in the triangulation of $\partial H$ is at least $3$. In other words, at most $3$ internal edges are incident to the same vertex of $\Omega$. We eliminate all internal edge patterns having fewer than $3$ internal edges incident to some vertex.
Now we consider the remaining internal edge patterns. For each vertex $(q)$ we have the list $\{(q,v_{i})\ |\ i=1,...,K_{q}\}$ of $K_{q}$ external edges incident to $q$. We write $K=K_{q}$ and order these $K=K_{q}$ vertices cyclically according to the link of $(q)$ in $\Omega$. The link of $L_{q}$ of $(q)$ in $\partial H$ is some permutation $(w_{1},...,w_{K})$ of $(v_{1},...,v_{K})$.
Lemma 2.5 (Cycle Rule)
In order to be a viable candidate for the link of $(q)$ in $\partial H$, the link $(w_{1},...,w_{K})$ must satisfy two properties.
  1. 1.
    $(w_{i},w_{i+1})$ must be an external edge for all $i$. Indices are taken cyclically
  2. 2.
    The cycle $(w_{1},...,w_{K})$ must be a dihedral permutation of $(v_{1},...,v_{K})$.
Proof: The necessity of Condition 1 is obvious. Condition 2 requires some explanation. Let $P$ be vertex $(q)$. Consider the picture in $\partial H$ at $P$. Since the points are in general position, $\partial H$ is a proper convex cone near $P$. We let $\Pi$ be a plane parallel to a support plane through $(q)$ that just cuts off a small corner of $H$ near $(q)$. Consider the intersection $\partial H\cap\Pi$. This is a convex $K$-gon $\Delta$, and the cyclic order of $\Delta$ is given by $(w_{1},...,w_{K})$.
Figure 2.1: $\Delta$ (black) and $\gamma$ (blue)
At the same time, and using the notation from the proof of Lemma 2.3, let $\gamma=\widehat{\Omega}_{P}\cap\Pi$. The polygonal loop $\gamma$ is contained in the region bounded by $\Delta$ and visits the vertices of $\Delta$ in the order $(v_{1},...,v_{K})$. Figure 2.1 shows some examples, with $\gamma$ slightly moved off $\Delta$ to make the drawing more clear. Note that if $K<6$ then there will be some extra vertices to $\gamma$ as well; this does not matter in the argument. If the permutation is not dihedral then $\gamma$ cannot be an embedded loop. But then $\Omega_{P}$ is not embedded, a contradiction. $\spadesuit$
Using the Cycle Rule (and our computer code) we eliminate all remaining internal edge patterns except $6$. These $6$ are the same up to combinatorial isomorphism. Figure 2.2 shows one of them. Evidently, all $6$ triangles incident to vertex $(2)$ are external triangles. This completes the proof of the Hull Lemma.
Figure 2.2: The one remaining pattern
2.4 The Hull Theorem
For the interested reader we sketch the proof of the result from [BE] that a $7$-vertex embedded polyhedral torus cannot have all $7$ vertices on its convex hull boundary. We call this result the Hull Theorem. This result is not needed for any of our results.
If such an example $\Omega$ exists then up to combinatorial isomorphism its internal edge pattern must be as in Figure 2.2. We think of $\Omega$ as a subset of projective space $\mbox{\boldmath{$P$}}^{3}$. We can apply a projective transformation so that vertex $(2)$ moves to the point $[0:0:1:0]$ at infinity in $\mbox{\boldmath{$P$}}^{3}$. This point is “infinitely far away” along the $Z$-axis. We also can arrange that the (now) rays $(2j)$ start at $(j)$ and move downward (rather than upwards) along the $Z$-axis, limiting on $(2)$. We do all this by a projective transformation that maps $H-\{(2)\}$ into the affine patch $\mbox{\boldmath{$R$}}^{3}$.
Now, $H-\{(2)\}$ is still a convex subset in $\mbox{\boldmath{$R$}}^{3}$. It is like a hexagonal prism that has been truncated at one end. Here is the crucial observation: since $H$ is convex, the projection of the hexagon $(603541)$ into the $XY$-plane is a convex hexagon. We can further normalize so that the projections of $(15)$ and $(36)$ into the $XY$-plane are parallel line segments. We do this by mapping the line $215]\cap[236]$ to a line at infinity which contains $(2)$. Here $[abc]$ is the plane containing $(a)$ and $(b)$ and $(c)$.
Figure 2.3: Projection of part of $\partial H$ into the $XY$-plane
The only triangles of $\partial H$ not in $\Omega$ are $(036)$ and $(145)$. By convexity, these two triangles are bent downward. What we mean is that the plane containing $(036)$ has the rest of $H$ beneath it. The same goes for the plane containing $(145)$.
The blue triangle in Figure 2.3 is the internal triangle $(346)$. We can foliate this triangle by parallel line segments as shown in the figure. These segments are parallel in space and they project to parallel segments in $\mbox{\boldmath{$R$}}^{2}$. The red triangle in Figure 2.3 is the internal triangle $(015)$. We make all the same constructions for this triangle.
Say that a special planes $\mbox{\boldmath{$R$}}^{3}$ whose projections to $\mbox{\boldmath{$R$}}^{2}$ is a line parallel to the projections of our red and blue foliations. One special plane contains $(15)$. In this plane, the red foliation is above the blue foliation. Another special plane contains $(36)$. In this plane, the blue foliation is above the red foliation. So, by the intermediate value theorem, there is a special plane for which these two foliations coincide. But then $(015)$ and $(346)$ intersect. This contradicts the fact that $\Omega$ is embedded.
3 Robust Embedding
3.1 Main Result
Let $T=T(z_{0},z_{1},z_{2})$ be the example from Equation 1. Let $T(z^{\prime}_{0},z^{\prime}_{1},z^{\prime}_{2})$ be as in Equation 1, but with variables $z_{0}^{\prime},z_{1}^{\prime},z_{2}^{\prime}$. We do not insist that $z_{0}^{\prime},z_{1}^{\prime},z_{2}^{\prime}$ are integers. Let
$|T^{\prime}-T|=\max(|z_{0}^{\prime}-z_{0}|,|z_{1}^{\prime}-z_{1}|,|z_{2}^{\prime}-z_{2}|).$ (5)
$T$ is $K$-robustly embedded if and only if $T^{\prime}$ is embedded whenever $|T^{\prime}-T|\leq K$. Here is our main result.
Lemma 3.1 (Robust Embedding)
$T$ is $10^{-4}$ robustly embedded.
The rest of this section gives the proof. We have $120=24+72+24$ triangle pairs to consider. The first $24$ are pairs having 0 vertices in common. The next $72$ are pairs having $1$ vertex in common. The last $24$ are pairs having $2$ vertices in common. We don’t need to check these last $24$.
Lemma 3.2
Let $T^{\prime}$ be a perturbation of $T$ such that all triangle pairs with no vertices in common are disjoint and all pairs with one vertex in common only intersect at the common vertex. Then $T^{\prime}$ is embedded.
Proof: We will argue by contradiction. If $T^{\prime}$ is not embedded, then we can find a pair of triangles $\tau_{1},\tau_{2}$ having two vertices $v,w$ in common, which overlap on an open set. Let $x_{j}$ be the third vertex of $\tau_{j}$. After relabeling if necessary, we can assume that some initial segment $s$ of the ray $\overrightarrow{vx_{1}}$ lies in $\tau_{2}$. But then let $\beta$ be the triangle of $T^{\prime}$ that shares the edge $vx_{1}$ with $\tau_{1}$. The triangle $\beta$ only has the vertex $v$ in common with $\tau_{2}$ but nevertheless intersects $\tau_{2}$ in a set that contains the segment $s$. This is a contradiction. $\spadesuit$
Suppose $(\Delta_{0},\Delta_{1})$ is a pair of triangles with no common vertices. Given a vector $L$ we define
$m_{j}(L)=\min_{v\in\Delta_{j}}v\cdot L,\hskip 30.0ptM_{j}(L)=\max_{v\in\Delta_{j}}v\cdot L.$ (6)
By convexity, we can compute these quantities just by computing the dot products on the vertices.
We call $(\Delta_{0},\Delta_{1})$ $\lambda$-separated if there exists a vector $L$ with $\|L\|_{\infty}\leq 1$ such that one of the two equations holds:
$M_{0}(L)+\lambda<m_{1}(L),\hskip 30.0ptM_{1}(L)+\lambda<m_{0}(L),$ (7)
The quantity $\|\cdot\|_{\infty}$ is the maximum of the absolute values of the coordinates. We have $\Delta_{0}\cap\Delta_{1}=\emptyset$ as long as this pair is 0-separated.
We show by direct calculation that all $24$ pairs are $10^{-3}$ separated. This is an integer calculation. For the purposes of calculation, we scale $T$ up by $10^{32}$. The scaled polyhedral torus $\Omega$ has integer vertices. (This is just for calculational purposes; our results are about $T$ itself.) Then we exhibit integer vectors $L_{0},...,L_{23}$ with $\|L\|_{\infty}\leq 10^{6}$ such that Equation 7 holds, with respect to the triangle pairs in $\Omega$, for $\lambda=10^{35}$. When we scale everything back down we get our $10^{-3}$-separation. Here are the triangle pairs and the corresponding vectors.
$\begin{matrix}\begin{bmatrix}0\\ 1\\ 2\end{bmatrix}\ \begin{bmatrix}3\\ 4\\ 6\end{bmatrix}&\begin{bmatrix}0\\ 1\\ 2\end{bmatrix}\ \begin{bmatrix}3\\ 5\\ 7\end{bmatrix}&\begin{bmatrix}0\\ 1\\ 2\end{bmatrix}\ \begin{bmatrix}3\\ 6\\ 7\end{bmatrix}&\begin{bmatrix}0\\ 1\\ 3\end{bmatrix}\ \begin{bmatrix}2\\ 4\\ 7\end{bmatrix}&\begin{bmatrix}0\\ 1\\ 3\end{bmatrix}\ \begin{bmatrix}2\\ 5\\ 6\end{bmatrix}&\begin{bmatrix}0\\ 1\\ 3\end{bmatrix}\ \begin{bmatrix}2\\ 6\\ 7\end{bmatrix}&\begin{bmatrix}0\\ 2\\ 4\end{bmatrix}\ \begin{bmatrix}1\\ 5\\ 7\end{bmatrix}&\begin{bmatrix}0\\ 2\\ 4\end{bmatrix}\ \begin{bmatrix}3\\ 5\\ 7\end{bmatrix}\\[12.0pt] \begin{bmatrix}0\\ 2\\ 4\end{bmatrix}\ \begin{bmatrix}3\\ 6\\ 7\end{bmatrix}&\begin{bmatrix}0\\ 3\\ 5\end{bmatrix}\ \begin{bmatrix}1\\ 4\\ 7\end{bmatrix}&\begin{bmatrix}0\\ 3\\ 5\end{bmatrix}\ \begin{bmatrix}2\\ 4\\ 7\end{bmatrix}&\begin{bmatrix}0\\ 3\\ 5\end{bmatrix}\ \begin{bmatrix}2\\ 6\\ 7\end{bmatrix}&\begin{bmatrix}0\\ 4\\ 6\end{bmatrix}\ \begin{bmatrix}1\\ 2\\ 5\end{bmatrix}&\begin{bmatrix}0\\ 4\\ 6\end{bmatrix}\ \begin{bmatrix}1\\ 5\\ 7\end{bmatrix}&\begin{bmatrix}0\\ 4\\ 6\end{bmatrix}\ \begin{bmatrix}3\\ 5\\ 7\end{bmatrix}&\begin{bmatrix}0\\ 5\\ 6\end{bmatrix}\ \begin{bmatrix}1\\ 3\\ 4\end{bmatrix}\\[12.0pt] \begin{bmatrix}0\\ 5\\ 6\end{bmatrix}\ \begin{bmatrix}1\\ 4\\ 7\end{bmatrix}&\begin{bmatrix}0\\ 5\\ 6\end{bmatrix}\ \begin{bmatrix}2\\ 4\\ 7\end{bmatrix}&\begin{bmatrix}1\\ 2\\ 5\end{bmatrix}\ \begin{bmatrix}3\\ 4\\ 6\end{bmatrix}&\begin{bmatrix}1\\ 2\\ 5\end{bmatrix}\ \begin{bmatrix}3\\ 6\\ 7\end{bmatrix}&\begin{bmatrix}1\\ 3\\ 4\end{bmatrix}\ \begin{bmatrix}2\\ 5\\ 6\end{bmatrix}&\begin{bmatrix}1\\ 3\\ 4\end{bmatrix}\ \begin{bmatrix}2\\ 6\\ 7\end{bmatrix}&\begin{bmatrix}1\\ 4\\ 7\end{bmatrix}\ \begin{bmatrix}2\\ 5\\ 6\end{bmatrix}&\begin{bmatrix}1\\ 5\\ 7\end{bmatrix}\ \begin{bmatrix}3\\ 4\\ 6\end{bmatrix}\end{matrix}$
$\begin{matrix}\begin{bmatrix}237139\\ -641458\\ 729586\end{bmatrix}&\begin{bmatrix}-246370\\ -672576\\ -697815\end{bmatrix}&\begin{bmatrix}-246370\\ -672576\\ -697815\end{bmatrix}&\begin{bmatrix}177844\\ -57056\\ 982403\end{bmatrix}&\begin{bmatrix}-14224\\ -292\\ -999899\end{bmatrix}&\begin{bmatrix}-14224\\ -292\\ -999899\end{bmatrix}&\begin{bmatrix}-708489\\ 225606\\ -668690\end{bmatrix}&\begin{bmatrix}333143\\ -637360\\ -694831\end{bmatrix}\\[12.0pt] \begin{bmatrix}-701892\\ 659986\\ -267892\end{bmatrix}&\begin{bmatrix}844211\\ 423426\\ 328660\end{bmatrix}&\begin{bmatrix}936233\\ 284018\\ -206884\end{bmatrix}&\begin{bmatrix}936233\\ 284018\\ -206884\end{bmatrix}&\begin{bmatrix}237139\\ -641458\\ 729586\end{bmatrix}&\begin{bmatrix}237139\\ -641458\\ 729586\end{bmatrix}&\begin{bmatrix}695629\\ -221591\\ -683372\end{bmatrix}&\begin{bmatrix}-24968\\ 3173\\ 999683\end{bmatrix}\\[12.0pt] \begin{bmatrix}-926243\\ -145342\\ -347781\end{bmatrix}&\begin{bmatrix}-990121\\ -140214\\ -957\end{bmatrix}&\begin{bmatrix}237139\\ -641458\\ 729586\end{bmatrix}&\begin{bmatrix}237139\\ -641458\\ 729586\end{bmatrix}&\begin{bmatrix}-7215\\ -6322\\ 999953\end{bmatrix}&\begin{bmatrix}-7215\\ -6322\\ 999953\end{bmatrix}&\begin{bmatrix}-926243\\ -145342\\ -347781\end{bmatrix}&\begin{bmatrix}237139\\ -641458\\ 729586\end{bmatrix}\end{matrix}$
Let $T^{\prime}$ be another example with $|T^{\prime}-T|\leq 10^{-4}$. Consider the triangle pair $(\Delta_{0}^{\prime},\Delta_{1}^{\prime})$ of $T^{\prime}$ that corresponds to $(\Delta_{0},\Delta_{1})$. In moving from $\Delta_{j}$ to $\Delta_{j}^{\prime}$, no vertex moves by more than $10^{-4}$, and at most one of the coordinates changes per vertex. Hence, each dot product above changes by at most $10^{-4}$. By the triangle inequality, $(\Delta_{0}^{\prime},\Delta_{1}^{\prime})$ is 0-separated. Hence $\Delta_{0}^{\prime}\cap\Delta_{1}^{\prime}=\emptyset$.
Now we consider when $(\Delta_{0},\Delta_{1})$ have a single common vertex $v$. Let $\Delta_{j}^{\sharp}$ be the edge of $\Delta_{j}$ opposite $v$. Define
$m_{j}^{\sharp}(L)=\min_{w\in\Delta^{\sharp}_{j}}w\cdot L,\hskip 30.0ptM_{j}^{\sharp}(L)=\max_{w\in\Delta^{\sharp}_{j}}w\cdot L.$ (8)
This time we call $(\Delta_{0},\Delta_{1})$ $\lambda$-separated if there exists a vector $L$ with $\|L\|_{\infty}\leq 1$ such that one of the two equations holds:
$M^{\sharp}_{0}(L)+\lambda<v\cdot L<m^{\sharp}_{1}(L)-\lambda,\hskip 30.0ptM^{\sharp}_{1}(L)+\lambda<v\cdot L<m^{\sharp}_{0}(L)-\lambda,$ (9)
If $(\Delta_{0},\Delta_{1})$ is 0-separated then $\Delta_{0}\cap\Delta_{1}=\{v\}$. Using the same scaling trick, we show by direct calculation that all $72$ pairs are $10^{-3}$-separated. Here are the vectors.
$\begin{matrix}\begin{bmatrix}0\\ 1\\ 2\end{bmatrix}\ \begin{bmatrix}0\\ 3\\ 5\end{bmatrix}&\begin{bmatrix}0\\ 1\\ 2\end{bmatrix}\ \begin{bmatrix}0\\ 4\\ 6\end{bmatrix}&\begin{bmatrix}0\\ 1\\ 2\end{bmatrix}\ \begin{bmatrix}0\\ 5\\ 6\end{bmatrix}&\begin{bmatrix}0\\ 1\\ 3\end{bmatrix}\ \begin{bmatrix}0\\ 2\\ 4\end{bmatrix}&\begin{bmatrix}0\\ 1\\ 3\end{bmatrix}\ \begin{bmatrix}0\\ 4\\ 6\end{bmatrix}&\begin{bmatrix}0\\ 1\\ 3\end{bmatrix}\ \begin{bmatrix}0\\ 5\\ 6\end{bmatrix}&\begin{bmatrix}0\\ 2\\ 4\end{bmatrix}\ \begin{bmatrix}0\\ 3\\ 5\end{bmatrix}&\begin{bmatrix}0\\ 2\\ 4\end{bmatrix}\ \begin{bmatrix}0\\ 5\\ 6\end{bmatrix}\\[12.0pt] \begin{bmatrix}0\\ 3\\ 5\end{bmatrix}\ \begin{bmatrix}0\\ 4\\ 6\end{bmatrix}&\begin{bmatrix}1\\ 2\\ 0\end{bmatrix}\ \begin{bmatrix}1\\ 3\\ 4\end{bmatrix}&\begin{bmatrix}1\\ 2\\ 0\end{bmatrix}\ \begin{bmatrix}1\\ 4\\ 7\end{bmatrix}&\begin{bmatrix}1\\ 2\\ 0\end{bmatrix}\ \begin{bmatrix}1\\ 5\\ 7\end{bmatrix}&\begin{bmatrix}1\\ 3\\ 0\end{bmatrix}\ \begin{bmatrix}1\\ 2\\ 5\end{bmatrix}&\begin{bmatrix}1\\ 3\\ 0\end{bmatrix}\ \begin{bmatrix}1\\ 4\\ 7\end{bmatrix}&\begin{bmatrix}1\\ 3\\ 0\end{bmatrix}\ \begin{bmatrix}1\\ 5\\ 7\end{bmatrix}&\begin{bmatrix}1\\ 2\\ 5\end{bmatrix}\ \begin{bmatrix}1\\ 3\\ 4\end{bmatrix}\\[12.0pt] \begin{bmatrix}1\\ 2\\ 5\end{bmatrix}\ \begin{bmatrix}1\\ 4\\ 7\end{bmatrix}&\begin{bmatrix}1\\ 3\\ 4\end{bmatrix}\ \begin{bmatrix}1\\ 5\\ 7\end{bmatrix}&\begin{bmatrix}2\\ 0\\ 1\end{bmatrix}\ \begin{bmatrix}2\\ 4\\ 7\end{bmatrix}&\begin{bmatrix}2\\ 0\\ 1\end{bmatrix}\ \begin{bmatrix}2\\ 5\\ 6\end{bmatrix}&\begin{bmatrix}2\\ 0\\ 1\end{bmatrix}\ \begin{bmatrix}2\\ 6\\ 7\end{bmatrix}&\begin{bmatrix}2\\ 4\\ 0\end{bmatrix}\ \begin{bmatrix}2\\ 5\\ 1\end{bmatrix}&\begin{bmatrix}2\\ 4\\ 0\end{bmatrix}\ \begin{bmatrix}2\\ 5\\ 6\end{bmatrix}&\begin{bmatrix}2\\ 4\\ 0\end{bmatrix}\ \begin{bmatrix}2\\ 6\\ 7\end{bmatrix}\\[12.0pt] \begin{bmatrix}2\\ 5\\ 1\end{bmatrix}\ \begin{bmatrix}2\\ 4\\ 7\end{bmatrix}&\begin{bmatrix}2\\ 5\\ 1\end{bmatrix}\ \begin{bmatrix}2\\ 6\\ 7\end{bmatrix}&\begin{bmatrix}2\\ 4\\ 7\end{bmatrix}\ \begin{bmatrix}2\\ 5\\ 6\end{bmatrix}&\begin{bmatrix}3\\ 0\\ 1\end{bmatrix}\ \begin{bmatrix}3\\ 4\\ 6\end{bmatrix}&\begin{bmatrix}3\\ 0\\ 1\end{bmatrix}\ \begin{bmatrix}3\\ 5\\ 7\end{bmatrix}&\begin{bmatrix}3\\ 0\\ 1\end{bmatrix}\ \begin{bmatrix}3\\ 6\\ 7\end{bmatrix}&\begin{bmatrix}3\\ 5\\ 0\end{bmatrix}\ \begin{bmatrix}3\\ 4\\ 1\end{bmatrix}&\begin{bmatrix}3\\ 5\\ 0\end{bmatrix}\ \begin{bmatrix}3\\ 4\\ 6\end{bmatrix}\\[12.0pt] \begin{bmatrix}3\\ 5\\ 0\end{bmatrix}\ \begin{bmatrix}3\\ 6\\ 7\end{bmatrix}&\begin{bmatrix}3\\ 4\\ 1\end{bmatrix}\ \begin{bmatrix}3\\ 5\\ 7\end{bmatrix}&\begin{bmatrix}3\\ 4\\ 1\end{bmatrix}\ \begin{bmatrix}3\\ 6\\ 7\end{bmatrix}&\begin{bmatrix}3\\ 4\\ 6\end{bmatrix}\ \begin{bmatrix}3\\ 5\\ 7\end{bmatrix}&\begin{bmatrix}4\\ 0\\ 2\end{bmatrix}\ \begin{bmatrix}4\\ 1\\ 3\end{bmatrix}&\begin{bmatrix}4\\ 0\\ 2\end{bmatrix}\ \begin{bmatrix}4\\ 7\\ 1\end{bmatrix}&\begin{bmatrix}4\\ 0\\ 2\end{bmatrix}\ \begin{bmatrix}4\\ 6\\ 3\end{bmatrix}&\begin{bmatrix}4\\ 6\\ 0\end{bmatrix}\ \begin{bmatrix}4\\ 1\\ 3\end{bmatrix}\\[12.0pt] \begin{bmatrix}4\\ 6\\ 0\end{bmatrix}\ \begin{bmatrix}4\\ 7\\ 1\end{bmatrix}&\begin{bmatrix}4\\ 6\\ 0\end{bmatrix}\ \begin{bmatrix}4\\ 7\\ 2\end{bmatrix}&\begin{bmatrix}4\\ 1\\ 3\end{bmatrix}\ \begin{bmatrix}4\\ 7\\ 2\end{bmatrix}&\begin{bmatrix}4\\ 7\\ 1\end{bmatrix}\ \begin{bmatrix}4\\ 6\\ 3\end{bmatrix}&\begin{bmatrix}4\\ 7\\ 2\end{bmatrix}\ \begin{bmatrix}4\\ 6\\ 3\end{bmatrix}&\begin{bmatrix}5\\ 0\\ 3\end{bmatrix}\ \begin{bmatrix}5\\ 1\\ 2\end{bmatrix}&\begin{bmatrix}5\\ 0\\ 3\end{bmatrix}\ \begin{bmatrix}5\\ 7\\ 1\end{bmatrix}&\begin{bmatrix}5\\ 0\\ 3\end{bmatrix}\ \begin{bmatrix}5\\ 6\\ 2\end{bmatrix}\\[12.0pt] \begin{bmatrix}5\\ 6\\ 0\end{bmatrix}\ \begin{bmatrix}5\\ 1\\ 2\end{bmatrix}&\begin{bmatrix}5\\ 6\\ 0\end{bmatrix}\ \begin{bmatrix}5\\ 7\\ 1\end{bmatrix}&\begin{bmatrix}5\\ 6\\ 0\end{bmatrix}\ \begin{bmatrix}5\\ 7\\ 3\end{bmatrix}&\begin{bmatrix}5\\ 1\\ 2\end{bmatrix}\ \begin{bmatrix}5\\ 7\\ 3\end{bmatrix}&\begin{bmatrix}5\\ 7\\ 1\end{bmatrix}\ \begin{bmatrix}5\\ 6\\ 2\end{bmatrix}&\begin{bmatrix}5\\ 6\\ 2\end{bmatrix}\ \begin{bmatrix}5\\ 7\\ 3\end{bmatrix}&\begin{bmatrix}6\\ 0\\ 4\end{bmatrix}\ \begin{bmatrix}6\\ 2\\ 5\end{bmatrix}&\begin{bmatrix}6\\ 0\\ 4\end{bmatrix}\ \begin{bmatrix}6\\ 7\\ 2\end{bmatrix}\\[12.0pt] \begin{bmatrix}6\\ 0\\ 4\end{bmatrix}\ \begin{bmatrix}6\\ 7\\ 3\end{bmatrix}&\begin{bmatrix}6\\ 0\\ 5\end{bmatrix}\ \begin{bmatrix}6\\ 7\\ 2\end{bmatrix}&\begin{bmatrix}6\\ 0\\ 5\end{bmatrix}\ \begin{bmatrix}6\\ 3\\ 4\end{bmatrix}&\begin{bmatrix}6\\ 0\\ 5\end{bmatrix}\ \begin{bmatrix}6\\ 7\\ 3\end{bmatrix}&\begin{bmatrix}6\\ 2\\ 5\end{bmatrix}\ \begin{bmatrix}6\\ 3\\ 4\end{bmatrix}&\begin{bmatrix}6\\ 2\\ 5\end{bmatrix}\ \begin{bmatrix}6\\ 7\\ 3\end{bmatrix}&\begin{bmatrix}6\\ 7\\ 2\end{bmatrix}\ \begin{bmatrix}6\\ 3\\ 4\end{bmatrix}&\begin{bmatrix}7\\ 1\\ 4\end{bmatrix}\ \begin{bmatrix}7\\ 2\\ 6\end{bmatrix}\\[12.0pt] \begin{bmatrix}7\\ 1\\ 4\end{bmatrix}\ \begin{bmatrix}7\\ 3\\ 5\end{bmatrix}&\begin{bmatrix}7\\ 1\\ 4\end{bmatrix}\ \begin{bmatrix}7\\ 3\\ 6\end{bmatrix}&\begin{bmatrix}7\\ 1\\ 5\end{bmatrix}\ \begin{bmatrix}7\\ 2\\ 4\end{bmatrix}&\begin{bmatrix}7\\ 1\\ 5\end{bmatrix}\ \begin{bmatrix}7\\ 2\\ 6\end{bmatrix}&\begin{bmatrix}7\\ 1\\ 5\end{bmatrix}\ \begin{bmatrix}7\\ 3\\ 6\end{bmatrix}&\begin{bmatrix}7\\ 2\\ 4\end{bmatrix}\ \begin{bmatrix}7\\ 3\\ 5\end{bmatrix}&\begin{bmatrix}7\\ 2\\ 4\end{bmatrix}\ \begin{bmatrix}7\\ 3\\ 6\end{bmatrix}&\begin{bmatrix}7\\ 2\\ 6\end{bmatrix}\ \begin{bmatrix}7\\ 3\\ 5\end{bmatrix}\end{matrix}$
$\begin{matrix}\begin{bmatrix}333143\\ -637360\\ -694831\end{bmatrix}&\begin{bmatrix}237139\\ -641458\\ 729586\end{bmatrix}&\begin{bmatrix}333143\\ -637360\\ -694831\end{bmatrix}&\begin{bmatrix}-6292\\ 6509\\ -999960\end{bmatrix}&\begin{bmatrix}-4005\\ -7564\\ -999964\end{bmatrix}&\begin{bmatrix}-35967\\ 13057\\ 999267\end{bmatrix}&\begin{bmatrix}634978\\ -415069\\ -651553\end{bmatrix}&\begin{bmatrix}634978\\ -415069\\ -651553\end{bmatrix}\\[12.0pt] \begin{bmatrix}695629\\ -221591\\ -683372\end{bmatrix}&\begin{bmatrix}-246370\\ -672576\\ -697814\end{bmatrix}&\begin{bmatrix}-246370\\ -672576\\ -697814\end{bmatrix}&\begin{bmatrix}-246370\\ -672576\\ -697814\end{bmatrix}&\begin{bmatrix}237139\\ -641458\\ 729586\end{bmatrix}&\begin{bmatrix}844211\\ 423426\\ 328660\end{bmatrix}&\begin{bmatrix}-732046\\ -442762\\ -517756\end{bmatrix}&\begin{bmatrix}237139\\ -641458\\ 729586\end{bmatrix}\\[12.0pt] \begin{bmatrix}237139\\ -641458\\ 729586\end{bmatrix}&\begin{bmatrix}-385755\\ 246693\\ -889009\end{bmatrix}&\begin{bmatrix}166403\\ -44834\\ 985038\end{bmatrix}&\begin{bmatrix}-4495\\ -9329\\ -999947\end{bmatrix}&\begin{bmatrix}-4495\\ -9329\\ -999947\end{bmatrix}&\begin{bmatrix}-567238\\ 186300\\ -802206\end{bmatrix}&\begin{bmatrix}1668\\ -778\\ -999999\end{bmatrix}&\begin{bmatrix}1668\\ -778\\ -999999\end{bmatrix}\\[12.0pt] \begin{bmatrix}892975\\ -311517\\ 324889\end{bmatrix}&\begin{bmatrix}866799\\ -259538\\ 425793\end{bmatrix}&\begin{bmatrix}989545\\ 144137\\ 5078\end{bmatrix}&\begin{bmatrix}-246370\\ -672576\\ -697814\end{bmatrix}&\begin{bmatrix}-246370\\ -672576\\ -697814\end{bmatrix}&\begin{bmatrix}-246370\\ -672576\\ -697814\end{bmatrix}&\begin{bmatrix}572750\\ 676209\\ -463356\end{bmatrix}&\begin{bmatrix}572750\\ 676209\\ -463356\end{bmatrix}\\[12.0pt] \begin{bmatrix}107339\\ 785395\\ 609617\end{bmatrix}&\begin{bmatrix}572750\\ 676209\\ -463356\end{bmatrix}&\begin{bmatrix}333143\\ -637360\\ -694831\end{bmatrix}&\begin{bmatrix}572750\\ 676209\\ -463356\end{bmatrix}&\begin{bmatrix}4434\\ -9487\\ 999945\end{bmatrix}&\begin{bmatrix}-709108\\ 221839\\ -669293\end{bmatrix}&\begin{bmatrix}237139\\ -641458\\ 729586\end{bmatrix}&\begin{bmatrix}1668\\ -778\\ -999999\end{bmatrix}\\[12.0pt] \begin{bmatrix}-385755\\ 246693\\ -889009\end{bmatrix}&\begin{bmatrix}-385755\\ 246693\\ -889009\end{bmatrix}&\begin{bmatrix}177845\\ -57056\\ 982403\end{bmatrix}&\begin{bmatrix}237139\\ -641458\\ 729586\end{bmatrix}&\begin{bmatrix}237139\\ -641458\\ 729586\end{bmatrix}&\begin{bmatrix}-990563\\ 108272\\ 84040\end{bmatrix}&\begin{bmatrix}844211\\ 423426\\ 328660\end{bmatrix}&\begin{bmatrix}936233\\ 284018\\ -206884\end{bmatrix}\\[12.0pt] \begin{bmatrix}-990563\\ 108272\\ 84040\end{bmatrix}&\begin{bmatrix}-926242\\ -145342\\ -347781\end{bmatrix}&\begin{bmatrix}734445\\ -232226\\ -637701\end{bmatrix}&\begin{bmatrix}-990563\\ 108272\\ 84040\end{bmatrix}&\begin{bmatrix}-926242\\ -145342\\ -347781\end{bmatrix}&\begin{bmatrix}924375\\ 79788\\ -373049\end{bmatrix}&\begin{bmatrix}1108\\ -1139\\ 999999\end{bmatrix}&\begin{bmatrix}1108\\ -1139\\ 999999\end{bmatrix}\\[12.0pt] \begin{bmatrix}-749170\\ 250875\\ 613030\end{bmatrix}&\begin{bmatrix}-989884\\ -141568\\ 9366\end{bmatrix}&\begin{bmatrix}-14176\\ -6521\\ 999878\end{bmatrix}&\begin{bmatrix}-749170\\ 250875\\ 613030\end{bmatrix}&\begin{bmatrix}-14176\\ -6521\\ 999878\end{bmatrix}&\begin{bmatrix}-749170\\ 250875\\ 613030\end{bmatrix}&\begin{bmatrix}-14176\\ -6521\\ 999878\end{bmatrix}&\begin{bmatrix}-926242\\ -145342\\ -347781\end{bmatrix}\\[12.0pt] \begin{bmatrix}-926242\\ -145342\\ -347781\end{bmatrix}&\begin{bmatrix}-990563\\ 108272\\ 84040\end{bmatrix}&\begin{bmatrix}-734159\\ 232033\\ -638100\end{bmatrix}&\begin{bmatrix}965929\\ 52857\\ 253351\end{bmatrix}&\begin{bmatrix}237139\\ -641458\\ 729586\end{bmatrix}&\begin{bmatrix}-923184\\ -181427\\ 338845\end{bmatrix}&\begin{bmatrix}-990563\\ 108272\\ 84040\end{bmatrix}&\begin{bmatrix}-923184\\ -181427\\ 338845\end{bmatrix}\end{matrix}$
But now the same argument as above shows that for $T^{\prime}$ (as above) all such pairs are 0-separated. Hence $\Delta_{0}^{\prime}\cap\Delta_{1}^{\prime}=\{v^{\prime}\}$.
The two cases above combine with Lemma 3.2 to prove that $T^{\prime}$ is embedded when $|T^{\prime}-T|<10^{-4}$. Hence $T$ is $10^{-4}$-robustly embedded.
3.2 The Vectors
Logically speaking, it does not matter how we produced the vectors above, but here we describe our method. Define
$\phi(x)=2x^{\prime}-1,\qquad x^{\prime}=x-{\rm floor\/}(x).$ (10)
The range of $\phi$ is [1,1][-1,1]. Next, we define
$V_{n}=\frac{W_{n}}{\|W_{n}\|},\hskip 30.0ptW_{n}=(\phi(n\sqrt{2}),\phi(n\sqrt{3}),\phi(n\sqrt{5})).$ (11)
The countable collection $\{V_{n}\}$ is dense in the unit sphere. For each pair $(\Delta_{0},\Delta_{1})$ of $T$ we try to verify a numerical version of Equation 7 or 9 (as the case may be). We found by trial and error that taking $\lambda=5\times 10^{-3}$ in the no-common-vertex case and $\lambda=10^{-3}$ in the $1$-common vertex case works well. For each pair, the search yielded some $V_{n}$. We then define
$L_{n}={\rm floor\/}(10^{6}V_{n}),$ (12)
This operation is applied componentwise. This gives us our lists above.
4 Near Flatness
4.1 The Main Result
Let $T$ be the torus from Equation 1. We say that $T$ is $\epsilon$-flat if $\max_{a}|\theta_{a}-2\pi|<\epsilon.$ The maximum is taken over all $8$ cone angles. In this chapter we prove the following result using rational arithmetic.
Lemma 4.1 (Near Flat)
$T$ is $10^{-30}$-flat.
Let $\Omega$ be the integer torus we get by scaling $T$ up by $10^{32}$. It suffices to prove that $\Omega$ is $10^{-30}$-flat. We mention that we computed, using $100$ digit precision arithmetic and $100$ digit precision versions of the square-root and arccos functions, that $\Omega$ is $(7.26\times 10^{-32})$-flat. Some readers might accept this calculation as “good enough proof” and skip the rest of this chapter.
Let $\theta_{i}$ denote the cone angle of $\Omega$ at the $(i)$th vertex. Here we discuss the idea behind the rational arithmetic proof of the Near Flat Lemma. Figure 1.3 shows a plot of the intrinsic flat structure on $\Omega$ up to an error of about $10^{-16}$. Around each vertex $V_{i}$ we have $6$ triangles $T_{i0},...,T_{i5}$. These triangles do not necessarily fit perfectly in the plane, but the error is so small that you cannot tell from the naked eye. Call the union of these $6$ triangles a flower. The $(i)$th flower corresponds to the $(i)$th star in $\Omega$, though only approximately.
For our proof below, we compute high precision versions of the flowers, then scale the coordinates up by $10^{32}$, then round down all the coordinate to integers. This gives us integer models which approximate the geometry of the $8$ stars in $\Omega$. Now something very nice happens. For our scaled-up models, these big integer triangles fit together perfectly around a vertex. Put another way, for each flower $F$ there are $6$ integer vectors $V_{0},...,V_{5}$ such that each of the $6$ triangles in $F$ is defined by $2$ cyclically consecutive vectors on the list. We then compute the mismatch between the various dot products associated to $\Omega$ and the corresponding dot products associated to the integer flowers, and this gives us the Near Flat Lemma. We list the coordinates for our flowers at the end of the chapter.
Rather than compare the angles, we compare the dot products which go into the calculation of the angles. We finish off the argument by using the Lipschitz properties of the function that converts between the relevant dot product ratio and the angle. We perform exact integer arithmetic calculations using the BigInteger class. Out of a slight laziness, and a desire for a transparent argument over a tedious exercise that anyone can perform, we will avail ourselves of the BigDecimal class in Java, which is perfectly capable of computing the ratios of $64$ digit integers up to $100$ digits of precision.
4.2 The Proof
Let $\tau_{ab}$ be the $(b)$th triangle in the $(a)$th star of $\Omega$ and let $\tau_{ab}^{*}$ be the $(b)$th triangle in the $(a)$th integer flower. Let $V_{ab}$ and $W_{ab}$ be the two vectors that $\tau_{ab}$ defines. By this we mean that $V_{ab}$ points from vertex $a$ to the first other vertex of $\tau_{ab}$ and $W_{ab}$ points from vertex $a$ to the second other vertex of $\tau_{ab}$. (The torus is oriented, so we have no trouble defining “first” and “second” here.) We define $V_{ab}^{*}$ and $W_{ab}^{*}$ with respect to $\tau_{ab}^{*}$ in a similar manner. Let $\theta_{ab}$ and $\theta^{*}_{ab}$ respectively denote the angles of $\tau_{ab}$ and $\tau_{ab}^{*}$ at vertex $(a)$.
When $r>0$ we have
$\theta=\phi(r),\hskip 30.0pt\phi(\cdot)=\arccos(\sqrt{(\cdot)}),\qquad r=\frac{(V\cdot W)(V\cdot W)}{(V\cdot V)(W\cdot W)},$ (13)
We take the positive branch of the square root function here.
We have $48=8\times 6$ pairs of triangles to check. We get the following bounds which work uniformly for all these pairs:
  1. 1.
    $r,r^{*}\in I:=[.00112,.77714]$ for all $a,b$.
  2. 2.
    $|r-r^{*}|<1.07\times 10^{-32}$.
  3. 3.
    $V\cdot W>0$ and $V^{*}\cdot W^{*}>0$.
Lemma 4.2
$\phi$ is $15$-lipschitz on $I$.
Proof: We compute that
$\bigg{|}\frac{d\phi}{dt}\bigg{|}=\frac{1}{2\sqrt{t(1-t)}}.$
Establishing our bound is the same as establishing that
$225-\frac{1}{4t(1-t)}>0,\hskip 30.0pt\forall t\in I$
Establishing this bound is a routine exercise in algebra (or calculus) which we omit. $\spadesuit$
Combining Lemma 4.2 with our calculations, and remembering that our flowers have $6$ triangles in them, we see that $\Omega$ is $\epsilon$-flat when
$\epsilon=6\times 15\times 1.07\times 10^{-32}<10^{-30}.$
This completes the proof of the Near Flat Lemma.
4.3 The Integer Flowers
$\begin{bmatrix}124788644050109876602825361295674&0&0\\ 77712297261878079581593486900079&70129923825301614779354253224291&0\\ -70435981294870011898311359845207&186386084617464728269924163644524&0\\ -148148278556748091479904846745286&116256160792163113490569910420232&0\\ -98095839650912348417428186459704&-84460765042234115458262663828023&0\\ 53415963724939232100166950154753&-89151222668009434223254123690615&0\end{bmatrix}$
$\begin{bmatrix}84465310291672221527450441188266&0&0\\ 69550368811175622953657281341642&103609612884589392577024128236871&0\\ -34241316280724259280029981722340&108947382985381561571026048979621&0\\ -103791685091899882233687263063983&5337770100792168994001920742749&0\\ -88338793860668172702704004953884&-123183133805664642986337651698562&0\\ 44045801850979783812535215764631&-101107834979521394561120290119996&0\end{bmatrix}$
$\begin{bmatrix}104677635440723967179988723412878&0&0\\ 12034998257537687883214969911430&83603513440581754781321562053477&0\\ -94317081172076477932017348008492&54413409707656384418499405421507&0\\ -191630826036432582723537220778797&-128131286236400952463466737043786&0\\ -97313744864356104791519872770305&-182544695944057336881966142465293&0\\ 32097662001199218957367514938600&-185561707945141208999189687748482&0\end{bmatrix}$
$\begin{bmatrix}151584388889297911866446233050267&0&0\\ 56495441368939829023723230650622&94715849040541184436182819641071&0\\ -49223826337965360195502746428317&68639664650392174112843586146043&0\\ -124728889560435721214027036255644&-3861320096243261090662872882149&0\\ -68579908206882346371371696550992&-91317009315546418606880471182019&0\\ 56148981353553374842655339704651&-87455689219303157516217598299871&0\end{bmatrix}$
$\begin{bmatrix}199251097863976649156276288186992&0&0\\ 161120818626134715906111075736407&97485840852430004244786950552576&0\\ -44330167652533400164597159444753&121619286399136814961872050344316&0\\ -102277745880424119662590582022408&-18451785319703080875434679167189&0\\ -44113235471988020751569542690915&-116731435954750256516337414942322&0\\ 38130279237841933250165212450584&-97485840852430004244786950552577&0\end{bmatrix}$
$\begin{bmatrix}206863556376741568622946129706945&0&0\\ 3647481315941312863078263678069&108826590670126188269802629065616&0\\ -75789172382575842209671187858145&80117590481155307070115264260041&0\\ -116999860561305386939755847152992&-65757376990228984547591251604189&0\\ -41210688178729544730084659294848&-145874967471384291617706515864231&0\\ 58216001420386574195913238485630&-115617091154199741307099293537955&0\end{bmatrix}$
$\begin{bmatrix}188317304997339823174158679929020&0&0\\ 17841734520938330693404444524334&103145915434133734843279705318812&0\\ -80328865359490685097501135729666&26108658194654254413402678301333&0\\ -69693027059078945150953187346448&-83662494202455064363150318793564&0\\ 93593941525210768327841471388803&-210666025048595406864753841069280&0\\ 163286968584289713478794658735251&-127003530846140342501603522275717&0\end{bmatrix}$
$\begin{bmatrix}151584388889297911866446233050267&0&0\\ 95435407535744537023790893345615&87455689219303157516217598299870&0\\ -56239231360498057885401519560362&199072046789130364493850160222280&0\\ -105719267706905189219225977078940&-26076184390149010323339233495028&0\\ -56495441368939829023723230650623&-94715849040541184436182819641072&0\\ 95088947520358082842723002399644&-94715849040541184436182819641072&0\end{bmatrix}$
5 Proof of Theorem 1.2
5.1 A Topological Result
Throughout this chapter $\|\cdot\|$ denotes the Euclidean norm and $\|\cdot\|_{\infty}$ denotes the maximum of the absolute values of vector or matrix entries.
Let $G:\mbox{\boldmath{$R$}}^{3}\to\mbox{\boldmath{$R$}}^{3}$ be a smooth map. Let $B_{r}(p)$ denote the ball of radius $r$ about $p$. We say that $G$ is $\lambda$-expansive on $B_{r}(p)$ if, for all unit vectors $V$ and all $q\in B_{r}(p)$, we have the following conditions.
  • The angle between $dG_{p}(V)$ and $dG_{q}(V)$ is less than $\pi/3$.
  • $\|dG|_{q}(V)\|>2\lambda$.
Here we are interpreting $V$ as a vector in all tangent spaces simultaneously.
Lemma 5.1 (Surjectivity)
Let $q=G(p)$. If $G$ is $\lambda$-expansive on $B_{r}(p)$ then $B_{\lambda r}(q)\subset G(B_{r}(p))$.
Proof: Let $O$ be the origin and let $B=B_{1}(O)$. We can scale and translate so that $\lambda=1$ and $p=q=O$. We need to prove that $B\subset G(B)$.
Let $\alpha$ be any line segment connecting $O$ to a point on $\partial B$. Let $V_{\alpha}$ be the unit vector which coincides with $\alpha$. Our hypotheses on $G$ guarantees that the projection of $G(\alpha)$ onto $V_{\alpha}$ always moves with speed greater than $1$. Hence $G(\alpha\cap\partial B)$ lies outside $B$. Given the angle condition, the sphere $G(\partial B)$ is the generator of the homology group $H_{2}(\mbox{\boldmath{$R$}}^{3}-q)$ for any $q\in B$. This easily implies that $B\subset G(B)$. $\spadesuit$
5.2 The Proof Modulo a Bound
Let $F:(z_{0},z_{1},z_{2})=(\theta_{0},\theta_{1},\theta_{2})$ be the map from §1.5. Let $p\in\mbox{\boldmath{$R$}}^{3}$ be the point that corresponds to our torus from Equation 1.1. For the rest of the chapter let $B=B_{10^{-28}}(p)$. Below we will establish the following bound.
Lemma 5.2 (Crude Bound)
Throughout $B$, we have
$\bigg{|}\frac{\partial^{2}F_{k}}{\partial z_{i}\partial z_{j}}\bigg{|}<10^{25},\hskip 30.0pt\forall\ i,j,k\in\{0,1,2\}.$
Now we have all the ingredients needed for the proof of Theorem 1.2. Define
$M=\left[\begin{matrix}-0.91&+0.74&+0.39\\ +0.74&-1.92&+1.14\\ +0.39&+1.14&-0.06\end{matrix}\right]$ (14)
Lemma 5.3
$\|dF_{q}-M\|<1/50$ for all $q\in B$.
Proof: The equation for the Jacobian $dF_{p}$, given in Equation 3, tells us that
$\|M-dF_{p}\|_{\infty}<\frac{1}{100}.$ (15)
Integrating the bound from the Crude Bound Lemma along a rectilinear path of length at most $\sqrt{3}\times 10^{-28}$ that joins $p$ to $q$ we see that
$\|dF_{q}-dF_{p}\|_{\infty}<\sqrt{3}\times 10^{-28}\times 10^{25}<\frac{1}{100}.$
This combines with Equation 15 and the triangle inequality to prove this lemma. $\spadesuit$
Lemma 5.4 (Expansion)
$F$ is $(1/10)$-expansive on $B$.
Proof: Let $V$ be any unit vector. The eigenvalues of $M$ all exceed $0.76$ in absolute value. Since $M$ is a symmetric matrix, we have $\|M(V)\|\geq 3/4$.
By the previous lemma we can write $dF_{q}=M+E_{q}$ where $\|E_{q}\|_{\infty}<1/50$. Note that $\|E_{q}(V)\|<3/50$. Since $\|M(V)\|>3/4$ and $\|E_{q}(V)\|<3/50$, we have $\|E_{q}(V)\|/\|M(V)\|<2/25$, which easily implies that the angle between $M(V)$ and $M(V)+E_{q}(V)$ is less than $\pi/3$. We also have
$\|dF_{q}(V)\|\geq\|M(V)\|-\|E_{q}(V)\|\geq\frac{3}{4}-\frac{3}{50}>\frac{1}{5}.$
Hence $F$ is $(1/10)$-expansive in $B$. $\spadesuit$
Recall that our torus $T$ corresponds to the point $p$. Let $q=F(p)\in\mbox{\boldmath{$R$}}^{3}$ and also let $q^{*}=(2\pi,2\pi,2\pi)$. Since $T$ is $10^{-30}$-flat, we have $\|q-q^{*}\|_{\infty}<10^{-30}$. Hence $q^{*}\in B_{10^{-29}}(q)$. Combining this with the Expansion and Surjectivity Lemmas, we have
$q^{*}\in B_{10^{-29}}(q)\subset F(B_{10^{-28}}(p)).$ (16)
Hence there is some $p^{*}\in B_{10^{-28}}(p)$ such that $F(p^{*})=q^{*}$. But then the torus $T^{*}$ corresponding to $p^{*}$ is flat. Since $T$ is $10^{-4}$-robustly embedded, we see that $T^{*}$ is also embedded. This completes the proof of Theorem 1.2 modulo the Crude Bound Lemma.
5.3 Breaking Down the Terms
Before getting to the proof of the Crude Bound Lemma, I want to discuss a certain robustness in our argument because I don’t want our endgame to give the impression that somehow everything boils down to finely tuned estimates. We have arranged things so that we need a bound of $10^{25}$ on some partial derivatives. By working with more precision, we could have reduced the problem to needing a bound of, say, $10^{1000}$. Our $32$-digit example simultaneously gives us something that does not require us to sweat too much over the final estimate and something which lets us have relatively short data files.
Now we get to the details. Let $F(z_{0},z_{1},z_{2})=(\theta_{0},\theta_{1},\theta_{2})$ be the map we have been studying. We write the $(i,j)$th entry in the Jacobian $dF$ as $\partial_{i}\theta_{j}$. We are considering symmetric variations. We do not just move $(z_{0},z_{1},z_{2})$ but also we move $(z_{4},z_{3},z_{6})$ in a symmetric way. Thus, for instance
$\partial_{0}\theta_{j}=\frac{\partial\theta_{j}}{\partial z_{0}}+\frac{\partial\theta_{j}}{\partial z_{4}}.$
When we take second derivatives we get $4$-term expressions. For instance
$\partial_{01}\theta_{j}=\frac{\partial^{2}\theta_{j}}{\partial z_{0}z_{1}}+\frac{\partial^{2}\theta_{j}}{\partial z_{4}z_{1}}+\frac{\partial^{2}\theta_{j}}{\partial z_{0}z_{3}}+\frac{\partial^{2}\theta_{j}}{\partial z_{4}z_{3}}$
Each $\theta_{j}$ is a $6$-term expression of terms of the form $\theta_{ijk}$, where $\theta_{ijk}$ is the angle of triangle $(i,j,k)$ at vertex $i$. Hence, $\partial_{ij}\theta_{j}$ is a $24$-term sum of expressions of the form
$\partial(a,b,c,d,e)=\frac{\partial^{2}\theta_{cde}}{\partial z_{a}\partial z_{b}}.$ (17)
These expressions vanish when $\{a,b\}\not\subset\{c,d,e\}$.
We call $\partial(a,b,c,d,e)$ friendly if $c\not\in\{a,b\}$. Otherwise we call the term unfriendly. Our terminology indicates our preference.
Lemma 5.5
Every unfriendly term is a sum of at most $2$ friendly terms.
Proof: First consider terms of the form $\partial(a,a,a,b,c)$. Using the relation $\theta_{a}+\theta_{b}+\theta_{c}=\pi$ we have
$\partial(a,a,a,b,c)=-\partial(a,a,b,c,a)-\partial(a,a,c,a,b).$
The terms on the right are friendly.
Given that $\partial(a,b,c,d,e)=\partial(b,a,c,d,e)$, the only remaining unfriendly terms we have to treat are those of the form $\partial(b,a,a,b,c)$ and $\partial(b,a,a,c,b)$. We will treat the first one. The second one has the same treatment.
Thinking of our partial derivatives as operators on the function $\theta_{cde}$, we have the basic formula
$\frac{\partial}{\partial z_{c}}+\frac{\partial}{\partial z_{d}}+\frac{\partial}{\partial z_{e}}=0.$
Geometrically, the variation we are considering simply translates the triangle up or down along the $Z$-axis and changes none of the angles. Using this relation, we have
$\partial(b,a,a,b,c)=-\partial(b,b,a,b,c)-\partial(b,c,a,b,c).$
Both terms on the right are friendly. $\spadesuit$
In summary each partial derivative $\partial_{ij}\theta_{k}$ is a sum of $K$ friendly terms of the form $\partial(a,b,c,d,e)$ where $K\leq 24\times 2=48$. In view of this breakdown, it suffices to show $|\partial(a,b,c,d,e)|<10^{23}$ for all friendly terms.
5.4 The Final Estimates
We write vectors as $V=(u,v,w)$. We say that a pair $(V_{1},V_{2})$ of vectors is normalized if they have the form
$V_{1}=(0,v_{1},w_{1}),\hskip 30.0ptV_{2}=(u_{2},v_{2},w_{2}).$ (18)
We say that a torus $T^{\prime}$ is relevant if it corresponds to a point in the ball of radius $10^{-27}$ about the point describing our example $T$ in Equation 1. We say that a pair of normalized vectors is relevant if it is obtained from a triangle $(P_{0},P_{1},P_{2})$ in a relevant torus by the following procedure.
  • We set $V_{1}^{\prime}=P_{1}-P_{0}$ and $V_{2}^{\prime}=P_{2}-P_{0}$.
  • We rotate $(V_{1}^{\prime},V_{2}^{\prime})$ about the $Z$-axis to get a normalized pair $(V_{1},V_{2})$.
We make this construction for all permutations of these points. Thus each triangle in a relevant torus gives rise to $6$ pairs of relevant vectors.
Let $\vartheta(V_{1},V_{2})$ be the angle between $V_{1}$ and $V_{2}$. If we have an arbitrary pair of vectors we can rotate about the $Z$-axis so that they are normalized. This rotation does not change the geometry of the triangle defined by these vectors and it does not change the dependence of this geometry on the $Z$-coordinate. For this reason, each friendly term discussed in the previous section has the form
$\partial_{ij}\vartheta=\frac{\partial^{2}\vartheta}{\partial w_{i}\partial w_{j}},\hskip 30.0pt\vartheta=\arccos\bigg{(}\frac{V_{1}\cdot V_{2}}{\sqrt{(V_{1}\cdot V_{1})(V_{2}\cdot V_{2})}}\bigg{)}$ (19)
Let $g_{ij}=(\partial_{ij}\vartheta)^{2}$. This is a rational function which we compute explicitly below. Note also that $g_{12}=g_{21}$. So, to prove the Crude Bound Lemma, we just need to prove for all relevant pairs that
$g_{ij}<10^{46},\hskip 30.0pt(i,j)=(1,2),(1,1),(2,2).$
Lemma 5.6
For each relevant pair $(V_{1},V_{2})$ of normalized vectors (in the notation above) we have $\|V_{1}\|,\|V_{2}\|\in[0.8,2.4]$ and $|u_{2}|>0.016$.
Proof: The norm bounds combine a direct calculation for $T$ with the observation that the norms change by at most $10^{-28}$ when we move to some other point in $B$.
The bound on $u_{2}$ is more subtle. For each “raw pair” $(V_{1}^{\prime},V_{2}^{\prime})$ associated to our triangle we compute
$a(V_{1}^{\prime},V_{2}^{\prime})=|u_{1}^{\prime}v_{2}^{\prime}-u_{2}^{\prime}v_{1}^{\prime}|.$
This quantity measures the area of the parallelogram spanned by our vectors when we project them into the plane. In particular $a(V_{1}^{\prime},V_{2}^{\prime})=a(V_{1},V_{2})$ when $(V_{1},V_{2})$ is the normalized pair associated to $(V_{1}^{\prime},V_{2}^{\prime})$. We check directly that $|a(V_{1}^{\prime},V_{2}^{\prime})|>0.02$ for all relevant triangles. We also see from Equation 1 that the projection of each vector of $T$ into the plane has length less than $1.2$. Hence $|v_{1}|<1.2$. Combining everything, we have
$|u_{2}|=\frac{|u_{2}||v_{1}|}{|v_{1}|}=\frac{|a(V_{1},V_{2})|}{|v_{1}|}=\frac{|a(V_{1}^{\prime},V_{2}^{\prime})|}{|v_{1}|}>\frac{0.02}{1.2}>0.016.$
This completes the proof. $\spadesuit$
To make our proof more robust, we will work with worse bounds:
$\|V_{1}\|^{2},\|V_{2}\|^{2}\in(10^{-1},10)\hskip 30.0pt|u_{2}|>10^{-2}.$ (20)
Case 1: We have
$g_{12}=\frac{u_{2}^{4}v_{1}^{4}}{\left(u_{2}^{2}(v_{1}^{2}+w_{1}^{2})+(v_{2}w_{1}-v_{1}w_{2})^{2}\right)^{3}}<\frac{10^{4}}{(10^{-4}\times 10^{-2})^{3}}<10^{46}.$ (21)
Case 2: We have
$g_{11}=\frac{v_{1}^{2}\left(-2w_{1}(v_{2}w_{1}-v_{1}w_{2})^{3}+u_{2}^{2}(v_{1}^{2}+w_{1}^{2})(v_{1}^{2}v_{2}-2v_{2}w_{1}^{2}+3v_{1}w_{1}w_{2})\right)^{2}}{(v_{1}^{2}+w_{1}^{2})^{4}\left(u_{2}^{2}(v_{1}^{2}+w_{1}^{2})+(v_{2}w_{1}-v_{1}w_{2})^{2}\right)^{3}}<^{*}$
$\frac{32\times 80\times 10^{16}}{(10^{-1})^{4}\times(10^{-4}\times 10^{-1})^{3}}<\frac{10^{20}}{10^{-19}}<10^{46}.$ (22)
The starred inequality comes from the fact that when we expand the numerator out completely it has $32$ terms, all coefficients less than or equal to $80$ in absolute value, and maximum degree $16$.
Case 3: We have
$g_{22}=\frac{\left(\begin{aligned} &2v_{2}w_{2}(v_{2}w_{1}-v_{1}w_{2})^{3}+u_{2}^{4}(v_{1}^{3}v_{2}+3v_{1}^{2}w_{1}w_{2}+2w_{1}^{3}w_{2})\\ &+u_{2}^{2}\big{(}4v_{2}^{2}w_{1}^{3}w_{2}-6v_{1}v_{2}w_{1}^{2}w_{2}^{2}+3v_{1}^{2}w_{1}w_{2}(v_{2}^{2}+w_{2}^{2})+v_{1}^{3}(v_{2}^{3}-v_{2}w_{2}^{2})\big{)}\end{aligned}\right)^{2}}{\left(u_{2}^{2}+v_{2}^{2}+w_{2}^{2}\right)^{4}\left(u_{2}^{2}(v_{1}^{2}+w_{1}^{2})+(v_{2}w_{1}-v_{1}w_{2})^{2}\right)^{3}}<^{*}$
$\frac{55\times 124\times 10^{16}}{(10^{-1})^{4}\times(10^{-4}\times 10^{-1})^{3}}<\frac{10^{20}}{10^{-19}}<10^{46}.$ (23)
The starred inequality has the same explanation as in Case 2.
This completes the proof of the Crude Bound Lemma, and thereby completes the proof of Theorem 1.2. An $8$-vertex paper torus exists!
6 Appendix: Coordinates and Pictures
6.1 Another Pup Tent
After writing the first version of this paper, I fooled around with the coordinates and found coordinates for a pup tent that are simpler than the ones in Equation 1. Here are the coordinates up to $32$ digits.
[…]
The way I found this example is that I had ChatGPT-5 rotate the original example around the $Z$-axis, looking for rotations which can be truncated to two decimal places with only small changes. So, for instance, perhaps the first coordinate was .7803.7803, and there was not much to change. I certainly could have done this without the help of ChatGPT, but I was feeling lazy.
The reason to have very simple coordinates, aside from the aesthetic value, would be to try to recognize the $z_{0},z_{1},z_{2}$ coordinates as algebraic numbers. This would be nice, because then one could have an explicit $8$-vertex paper torus rather than one defined by the Implicit Function Theorem. I have not had any success doing this, however.
6.2 Two More Pictures
The next picture we show is a layout for building your own pup tent. The figure gives a folding guide and the fold lines for a fundamental domain of the torus. The picture is decorated with one of the level set foliations. When the thing is correctly folded and suitably rotated, the level curves all lie in horizontal planes.
The final picture shows some $3D$ pictures that Peter Doyle made. The left and middle columns show two annuli that together form the torus in the right column. The rows correspond to viewing the tent from the top, bottom, broadside, and head on.